# ToDo (#1): #2

### v2

@@ 2.1 The Ion Source
The $\mathrm{H}_{2}^{+}$ beam to be injected into the cyclotron and accelerated to 60 MeV/amu will be produced in a filament-driven multicusp ion source which we will refer to as MIST-1. The technical details of this ion source are reported in Refs. [25, 26, 27, 28, 29] and are summarized here.
2.1.1 Design
[…]
2.1.1.1 Gas Input
Hydrogen gas is introduced via an inlet in the backplate of the source as indicated in Fig. 2.2. The amount of hydrogen entering the ion source is controlled using a mass flow controller (MKS Instruments Model GV50A) which has a 5 SCCM full range. This system interfaces with the PC control system which is documented in detail in Ref. [29].
2.1.1.2 Ionization by the Filament
After the hydrogen gas is introduced into the chamber, it is ionized to form a plasma using electrons released from a tungsten filament by thermionic emission. The filament is mounted on DI water-cooled power feedthroughs in the back plate (cf. Fig. 2.2). The filament is heated by an (8.4V, 300 A max) power supply.
In addition to this power supply which is used to heat the filament, a second power supply is connected to the filament and to the source chamber to maintain a potential difference of 80-140V between them in order to facilitate electron discharge. The source body, back plate, and front plate are all mutually electrically insulated from each other for added flexibility in the electric field shape. A schematic of the wiring of the ion source is illustrated in Fig. 2.2
[…]
Figure 2.1: A 3D CAD rendering of the MIST-1 ion source.
Figure 2.1: A 3D CAD rendering of the MIST-1 ion source.
Figure 2.2: Wiring schematic for the ion source. The red square region corresponds to the HV reference potential. Thick blue lines are data cables. Black lines are power cables.
2.1.1.3 Confinement & Extraction of the Plasma
[…]
The electrons are then shaped and accelerated out via an extraction system (cf. Fig. 2.2). The extraction system comprises five electrodes which shape and accelerate the ion beam. These include a plasma electrode (+15 kV), a puller electrode (-3 to -20 kV), and three electrostatic lenses – Lens 1 (+10 kV to +15 kV), lens 2 (-3 kV to -15 kV), and lens 3 (+10 kV to +15 kV). The RFQ entrance flange is held at ground potential. We simulated beam formation and shaping in the extraction system with the IBSimu package [30], which performs 3D simulations including space charge. Due to the computational resources needed to perform full 3D simulations in IBSimu, we adopted a hybrid IBSimu (for the plasma modeling)/WARP (for the beam transport) [31] model for the final design. We show example trajectories in Fig. 2.3.
Figure 2.3: Cross section of the extraction system showing the ion beam moving through the extraction system prior to entering the RFQ. Simulation of the extraction system performed by IBSimu and WARP software.
@@ 2.1.2.1 Goals & Measurement Methods
Characterization of the ion source is done using a short diagnostic beam line. This test beam line comprises two Faraday cups, a dipole magnet with separation slits, and a pair of Allison emittance scanners. The total beam current is measured using the first Faraday cup located posterior to the extraction system. The dipole magnet with separation slips separates the ions by mass and passes them to the second Faraday cup which, in conjunction with the Allison scanners, measures each species horizontal and vertical emittances.
@@ 2.1.2.2 Optimization of Ion Source Parameters
  • Filament size, shape, & position (Optimization not yet studied)
  • Permenant magnet type (Optimization not yet studied)
[…]
Filament discharge voltage optimization. Two series of filament discharge voltage optimizations have been performed differing in how discharge current is managed. In the first, discharge current is permitted to freely find a steady state using the same method as used in the flow optimization described above. In the second, the discharge current was fixed at 4 A.
Filament discharge current optimization. One series of current optimizations was explored in which the flow rate and discharge were kept constant while discharge current was varied from 2 to 8 A. No strong dependence on current was observed.
@@ 2.1.2.3 Latest Optimized Results
[…]
@@
With current optimizations and a 4mm aperture, the ion source is capable of delivering approximately 1.1 mA of H+${}_{2}^{+}$ with 76% purity. With an 8 mm aperture this corresponds 4.4 mA. This is a world leading result and is only about a factor of 2 short of target.
@@
The main bottleneck for increased beam current was the extraction system, which was designed for initial testing and commissioning. We have recently replaced it with the version presented in this chapter. Our high fidelity simulations suggest that the new extraction system will be able to handle the nominal beam current. Measurements are planned for summer 2024.
@@ 2.2 The Low Energy Beam Transport
Details of the LEBT design and performance were given in Refs. [32, 33, 34] here we summarize and update these.
@@ 2.2.1 Design Overview
At this stage, ions emerge from the ion source. The beam now passes through a gap surrounded by a series of six electrodes arranged axially within a six-way cross (see Fig. 2.2). These six electrodes can have their voltage optimized in order to best match the parameters for acceptance into the RFQ. A conical flange is connected to the end of the six-way cross. The conical shape minimizes beam loss on exit.
Following the six-way cross lies a diagnostic segment. A gate valve separates the six-way cross from this segment. The diagnostic section comprises a second cross, which can be used for additional vacuum attachments, diagnostic readouts, and a chopping electrode. The diagnostic section also includes a solenoid and two magnetic steerers which can be used for additional beam tuning. A flange in place toward the end allows space for an ACCT which characterizes the beam immediately prior to entering the RFQ. The diagnostic segment is illustrated in Fig. 2.3.
@@ 2.2.2 Beam Steering & Chopping
The diagnostic segment also serves to steer the beam into the RFQ or terminate the beam as needed. Even small, sub-millimeter misalignment can prevent efficient injection of the beam into the RFQ. In order to ensure that the beam enters the RFQ at both the correct position and angle, two magnetic steerers and a solenoid are placed around the beamline. These are controlled independently by an external power source to produce tunable dipole fields which optimize beam entry into the RFQ. The optimal fields were calculated using COMSOL and Warp. These are discussed further in the following section.
In order to control the amount of power delivered from the ion source, a beam chopper is built into the cross of the diagnostic segment to allow for a duty factor in an otherwise DC beam. The beam chopper is built into the cross in the diagnostic section. The chopper comprises a copper ring surrounded by a steel beampipe which in turn is surrounded by a copper water cooling jacket. The copper ring can be put at a high potential relative to the cross, which is grounded. If the potential is sufficiently high the beam can be diverted such that it terminates. See Fig. 2.5. This method permits control over the amount of power delivered to the RFQ without disrupting the plasma itself.
A 12 mA beam of 15 keV particles fully terminating on the beam stop corresponds to $\sim$180 W power delivered. We need to ensure that the dump is capable of dissipating this amount of power. The system was modeled assuming a water cooling line at room temperature. To be conservative, the simulation was run using four times the actual anticipated beam power, i.e. 720 W. The beam dump never rose above 360 K in our simulation, significantly below the melting point of copper. The COMSOL simulation is shown in Fig. 2.5
Figure 2.4: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid. If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring.
Figure 2.4: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid. If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring.
Figure 2.5: COMSOL simulation of the beam terminating on the copper aperture surrounded by the steel beampipe, with copper water cooling channel on the exterior.

### v3

@@ 2.1 The Ion Source
The $\mathrm{H}_{2}^{+}$ beam to be injected into the cyclotron and accelerated to 60 MeV/amu will be produced by a filament-driven multicusp ion source. We have built a first prototype, the MIST-1 (Multicusp Ion Source at MIT - version 1), the technical details and first measured results of which are reported in Refs. [axani_high_2015, winklehner_first_2018, winklehner_high-current_2021-1, winklehner_new_2022, weigel_epics_2023]. Here we summarize these results and present ongoing upgrade work (MIST-2). The upgrade to MIST-2 improves on the MIST-1 by implementing several lessons learned during the systematic measurements and runtime experiences described below. A publication on the MIST-2 design and performance is forthcoming.
2.1.1 Operating Principle
In a multicusp ion source [ehlers:multicusp1], a plasma is confined by a set of permanent magnets that are arranged around the plasma chamber with alternating polarity, thereby creating “cusps” in the magnetic field. On average, the electrons and ions produced in the center of the chamber see an increasing magnetic field as they travel outward, which eventually turns them back. One notable exception are the cusps, which are the natural loss points of this configuration.
If the source is filament-driven [^1] Other options exist, e.g., RF-driven., electrons are emitted from a hot filament. A plasma discharge is initiated and sustained by applying a potential difference between the filament and (parts of) the source chamber. Thus, the filament acts as the cathode, and the source chamber acts as the anode of a discharge. Electron-impact ionization with neutral molecules and atoms in the source chamber produces secondary electrons and ions.
Neutral gas is fed into the source to facilitate the generation of a high-density plasma. The type of gas determines the ion species produced in the source. Here we use H2 gas to form a plasma – and beam – comprising protons, $\mathrm{H}_{2}^{+}$, and H+${}_{3}^{+}$.
@@ 2.1.1 Operating Principle
[…]
2.1.2 Mechanical Design
Figure 2.1: Cross section of MIST-1 (left) with extraction system and LEBT. Also shown is an example set of ion trajectories simulated with IBSimu and WARP. From Ref. [waites_high_2022].
For visual reference as we discuss the details of the ion source, 3D CAD renderings of MIST-1 (with subsequent LEBT) and MIST-2 are provided in Figs. 2.1 and 2.2, respectively. The main ion source parameters are also summarized in Table 2.1.
The MIST-type ion sources comprise a plasma chamber surrounded by a “water-jacket,” a back plate, and a front plate. The water jacket lets us pump de-ionized (DI) water around the plasma chamber for cooling. The radial confinement magnets are mounted inside the water-jacket and are cooled directly. At 6.5 cm, the plasma chamber is short compared to other ion sources. This is a direct consequence of the short mean free path of $\mathrm{H}_{2}^{+}$ in the plasma/neutral gas of the ion source discussed above.
The nominal configuration for the confinement magnets is: Four rows of alternating magnets on the back plate, twelve alternating radial magnets, and a star formation of twelve smaller magnets on the front plate, also alternating in polarity. This leaves a small volume of low magnetic flux around the extraction aperture, allowing the ions to easily drift out of the source. All magnets are Sm2Co17. For MIST-2 we have begun studying different configurations (no magnets, Nd2Fe14B magnets). See below for more details.
The back plate holds four strips of confinement magnets, as well as the water-cooled filament feedthroughs, and the gas inlet (a KF NW-16 elbow). Temperature sensors (thermocouples) can be attached to the back plate to monitor the temperature of feedthroughs and magnets. The back plate itself is also water-cooled with cooling channels embedded in the plate. The filament is mounted to the high-current feedthroughs using molybdenum filament holders. The filament itself is made from tungsten and can have different thickness as well as shapes (this is a parameter currently being optimized).
The front plate holds the front confinement magnets as well as the plasma aperture, a 75%W-25%Cu (aka “Elkonite”) plate that is embedded in the front plate. We have several plasma apertures with different-size openings ranging from 4 mm to 8 mm to test beam quality and total extractable current for various aperture diameters.
As seen in Figs. 2.1 and 2.2, following the plasma aperture is the “extraction system” which shapes and guides the beam as it is initially formed. The MIST-1 extraction system comprises five electrodes which shape and accelerate the ion beam. These include a plasma electrode (+15 kV), a puller electrode (-3 to -20 kV), and three electrostatic lenses – Lens 1 (+10 kV to +15 kV), lens 2 (-3 kV to -15 kV), and lens 3 (+10 kV to +15 kV).
Figure 2.2: MIST-2 CAD rendering calling out the most important features.
2.1.2.1 MIST-2 vs. MIST-1
The changes we made in the MIST-2 design are mainly to improve cooling, assembly and maintenance, and diagnostics. They are:
  • Replace all stainless steel components with OFHC copper to improve heat transfer
  • Increase the cross-section areas of all water cooling channels
  • New high-power feedthroughs that are detachable from the back plate
  • An additional port on the back plate for a Langmuir probe
  • A redesigned extraction system. The new pentode extraction system comprises a water-cooled puller, an intermediate electrode, a negative electrode and a grounded final electrode. The more compact extraction system with negative potential followed by ground will help keep space charge compensation high in the LEBT and allows us to use a Faraday cup and emittance scanners for beam diagnostics directly after the source.
These changes will improve long-term running stability, reduce maintenance downtime, and allow for a better understanding of the plasma and beam conditions. Our high-fidelity simulations using IBSimu [kalvasIBSIMUThreedimensionalSimulation2010a] and WARP [vayNovelMethodsParticleInCell2012] suggest that the beam dynamics downstream of the six-way cross will not change between the MIST-1 and MIST-2 configurations. All particle simulations and designs still hold for LEBT, RFQ, and Cyclotron.
2.1.3 Electrical Design
Hydrogen gas is introduced via an inlet in the backplate of the source as indicated in Fig. 2.2. The amount of hydrogen entering the ion source is controlled using a mass flow controller (MKS Instruments Model GV50A) which has a 5 SCCM full range. This system interfaces with the EPICS PC control system which is documented in detail in Ref. [weigel_epics_2023].
The filament is heated by an TDK Lambda (8.4 V, 300 A max) power supply. In addition to this power supply which is used to heat the filament, a second power supply (Matsusada REK) is connected to the filament and to the source chamber to maintain a potential difference of 80 - 140 V between them in order to facilitate electron discharge of up to 24 A max (typical during our tests was 5-10 A). The source body, back plate, and front plate are all mutually electrically insulated from each other for added flexibility in the electric field shape. Power is transferred to the platform via two isolation transformers of 2 kVA each. A schematic of the wiring of the ion source is illustrated in Fig. 2.3.
The ion source (platform) is lifted to high voltage using a Matsusada AU 15 mA, 20 kV power supply, all lens elements in the extraction system are powered by Matsusada AU 7.5 mA, 20 kV power supplies of either negative or positive polarity.
The devices on the HV platform (control computer, filament discharge, filament heating power supply, etc.) are inside a Faraday cage on platform potential to reduce electromagnetic interference (EMF). Surrounding the platform and ion source is another Faraday cage at ground potential for personnel protection and EMF reduction.
Figure 2.3: Wiring schematic for the ion source. The red square region corresponds to the HV reference potential. Green lines are data connections. Black lines are power cables.
2.1.4 Beam Dynamics
We simulate beam formation and shaping in the extraction system with the IBSimu package [kalvasIBSIMUThreedimensionalSimulation2010a], which performs cylindrically symmetric 2D (aka “RZ”) and 3D simulations including space charge through an iterative process. IBSimu also contains a 1D plasma sheath model which is a good approximation for the plasma-beam interface at the source aperture. Our extraction simulations of the MIST-1 ion source followed by a WARP [vayNovelMethodsParticleInCell2012] simulation of the LEBT yielded excellent agreement with measurements [winklehner_first_2018, winklehner_new_2022]. Example trajectories of the new MIST-1 extraction system and RFQ matching LEBT are shown in Fig. 2.1.
An extraction system simulation of the MIST-2 system is shown in Fig. 2.4. The simulated normalized RMS emittance for the new MIST-2 extraction system using an 8 mm diameter plasma aperture is 0.077 $\pi$-mm-mrad. Simulated normalized RMS emittances for the MIST-1 extraction system were between 0.05 and 0.1 $\pi$-mm-mrad depending on source aperture and extracted current and agreed well with measurements [winklehner_first_2018].
Figure 2.4: IBSimu Simulation of MIST-2 extraction. Left: trajectories, Right: x-x’ phase space. Both MIST-1 and MIST-2 exhibit a very low emittance (see text).
@@ 2.1.5.1 Goals & Measurement Methods
Characterization of the ion source is done using a short diagnostic beam line. This test beam line comprises two Faraday cups, a dipole magnet with separation slits, and a pair of Allison emittance scanners. The total beam current is measured using the first Faraday cup located posterior to the extraction system. The dipole magnet separates the ions by mass and passes them to the second Faraday cup for mass analysis. Allison-type electrostatic emittance scanners measure each species’ horizontal and vertical emittances.
@@ 2.1.5.2 Optimization of Ion Source Parameters
  • Filament size, shape, & position (Optimization not yet fully studied)
  • Permanent magnet type
H2 input flow optimization. Three series of optimizations on the input gas flow rate have been performed. The three series differ in how the filament discharge voltage is managed. In the first series, the PID loop which stabilizes discharge was disabled allowing the discharge current to rise to the source steady-state value and then stabilized at that value by reengaging the PID. In the second series, the discharge current was kept fixed at 4 A. In the third series, the discharge current was reduced to keep the total extracted current fixed at 1.25 mA. Under each of these three conditions, the gas input rate was varied from 0.25 - 1.25 SCCM. This study demonstrates a general trend of greater H+${}_{2}^{+}$ production at higher discharge currents and, as theoretically expected, a higher fraction of H+${}_{3}^{+}$ with higher flow rates.
Discharge voltage optimization. Two series of discharge voltage optimizations have been performed differing in how discharge current is managed. In the first, discharge current is permitted to freely find a steady state using the same method as used in the flow optimization described above. In the second, the discharge current was fixed at 4 A.
Discharge current optimization. One series of current optimizations was explored in which the flow rate and discharge were kept constant while discharge current was varied from 2 to 8 A. No strong dependence on current was observed.
@@ 2.1.5.3 Latest Optimized Results
Highest total current density: 41 mA/cm2 with 31% H+${}_{2}^{+}$ purity.
Highest H+${}_{2}^{+}$ purity: 76% purity with 11.4 mA/cm2 total current density.
Highest H+${}_{2}^{+}$ current density with H+${}_{2}^{+}$ dominant: 10.4 mA/cm2 with 62% purity [^2] Note that this is the current density of $\mathrm{H}_{2}^{+}$ alone..
@@
With current optimizations and a 4mm aperture, the ion source is capable of delivering approximately 1.1 mA of H+${}_{2}^{+}$ with 76% purity. With an 8 mm aperture this corresponds to 4.4 mA. This is a world-leading result and is only about a factor of 2 short of the target.
@@
The main bottleneck for an increased beam current was the extraction system, which was designed for initial testing and commissioning. We have recently replaced it with the version presented in this chapter. Our high-fidelity simulations suggest that the new extraction system will be able to handle the nominal beam current. Measurements are planned for summer 2024.
2.1.5.4 Magnet Study
The choice of permanent magnet material and strength is one of the optimization parameters of this type of ion source. Here, plasma density, confinement time, and magnet thermal properties are important parameters to consider. Originally, we selected Sm2Co17 magnets for their higher Curie temperature. However, long-term running tests suggest that the magnet temperature is well controlled, which would allow the use of Nd2Fe14B magnets, if they yielded a better beam.
We aim to maximize the $\mathrm{H}_{2}^{+}$ fraction of the MIST-1 ion source. Within the plasma, atomic processes include ionization, dissociation, and recombination, leading to protons, $\mathrm{H}_{2}^{+}$, and H+${}_{3}^{+}$ in the beam. Magnets with a higher field strength yield higher plasma density and longer confinement of the ions in the plasma, which can alter the rates of these processes. Thus, we are interested in how different magnet types alter the ion fraction and total current. We investigated the current and ion fraction of the beam with no magnets, Sm2Co17 magnets (grade 28), and Nd2Fe14B (grade N50H). The configuration without magnets provides no confinement, and the Nd2Fe14B magnets provide the strongest confinement. We performed our measurements using the diagnostic beam line presented in Ref. [winklehner_first_2018], containing a dipole magnet and a Faraday cup with 8 mm entrance slit for mass analysis. We investigated each magnet type at a range of H2 mass flows (0.25 SCCM to 1.5 SCCM) and discharge voltages (60 V to 140 V).
In our preliminary analysis, we observed a trade-off between the total beam current and the $\mathrm{H}_{2}^{+}$ fraction: with lower confinement, the $\mathrm{H}_{2}^{+}$ fraction was high, but the current was low. This result was expected; lower confinement and lower plasma density leads to less time and fewer reactants for the ions to undergo processes that result in $\mathrm{H}_{2}^{+}$ dissociation or the pick-up of another proton (producing H+${}_{3}^{+}$). However, the lower plasma density also leads to a lower total beam current.
For MIST-2, the baseline design is now to use Nd2Fe14B magnets because of the higher current they provide. A paper on the detailed analysis of the magnet study is forthcoming.
2.1.5.5 Results with the MIST-2 back plate
Figure 2.5: A typical spectrum with the new MIST-2 back plate. The total beam current was 3.43 mA, the $\mathrm{H}_{2}^{+}$ fraction $\approx 60$%. For a discussion about the small contaminant peak, cf. text.
We already performed the first test of the MIST-2 components described in Section 2.1.2.1 by installing the new back plate (which is backward compatible) on MIST-1. We were able to run the source up to a stable beam current of 3.43 mA. We did this with the source set to 7.5 kV (the nominal source voltage of 15 kV requires a full beam current of 10 mA to form the appropriate plasma meniscus), puller set to -5 kV, lens 1 at 1.5 kV, lens 2 at -3 kV, lens 3 at 0.5 kV and the filament heating set to 150 A.
Due to the improved cooling and filament feedthroughs and holders, the source ran very stably; as we expected. A sample spectrum from this series of studies is shown in Fig. 2.5. Here, a small cluster of beam contaminants can still be seen, which we attribute to insufficient baking and out-gassing of source components, including Viton O-rings. Past experience has shown that vacuum baking the O-rings before installation (done here), overall good vacuum conditions, and baking the source and beam line before operation (not done in this instance), decreases the contaminants to a negligible level after a few hours of operation [winklehner_first_2018].
2.1.6 Risks and Mitigation
Risk: Ion Source Long-Time Stability. As IsoDAR is designed to run continuously for 5 years, the long-time stability is a major concern for the ion source. Hands-on maintenance of the ion source is going to be limited, because of the close proximity of the cyclotron.
Mitigation: Before the final technical design, we will develop alternative heating methods (external antenna and 2.45 GHz microwaves). MIST-1 and MIST-2 were being designed with an exchangeable backplate for quick filament replacement and easy further development of alternative heating methods. Another option is a dual-source operation using a switching magnet.
Risk: Ion source maximum $\mathrm{H}_{2}^{+}$ current is too low. The current densities seen from the MIST-1 ion source are still a factor 2 below the requirements for IsoDAR.
Mitigation: Our current upgrade path should lead us comfortably to a total extracted current of 12 mA, particularly considering that MIST-1 has demonstrated an exceptionally low emittance, which opens up the possibility of going to even larger extraction apertures (10 mm or 12 mm). However, we also have tested a 2.45 GHz flat-field ECR ion source (the VIS [alonso_isodar_2015]) in the past, from which we measured 15 mA of $\mathrm{H}_{2}^{+}$. The low ratio of 50% $\mathrm{H}_{2}^{+}$ (30 mA total current) deterred us, but as a mitigation, a VIS-type ion source paired with a dipole magnet is a viable alternative.
@@ 2.2 The Low Energy Beam Transport
Details of the LEBT design and performance were given in Refs. [waites_matching_2021, waites_low_2022, waites_high_2022] here we summarize and update these.
@@ 2.2.1 Design Overview
At this stage, ions emerge from the ion source. The beam now passes through a gap surrounded by a series of six electrodes arranged axially within a six-way cross (see Fig. 2.1). These six electrodes can have their voltage optimized in order to best match the parameters for acceptance into the rest of the LEBT. A conical flange is connected to the end of the six-way cross. The conical shape minimizes beam loss on exit.
Figure 2.6: Left: Photo of Alpha Magnetics, Inc. steerers [2562NotreDame]. Right: Result of LEBT steering simulations. It can be seen that several millimeters offset can be achieved without significant change in emittance. From Ref. [waites_high_2022].
Following the six-way cross lies a diagnostic segment. A gate valve separates the six-way cross from this segment. The diagnostic section comprises a second cross, which can be used for additional vacuum attachments, diagnostic readouts, and a chopping electrode. The diagnostic section also includes a solenoid and two magnetic steerers which can be used for additional beam tuning. A flange in place toward the end allows space for an AC Current Transformer (ACCT) [bergozACCT] which characterizes the beam immediately prior to entering the RFQ. The diagnostic segment is illustrated in Fig. 2.1.
@@ 2.2.2 Beam Steering & Chopping
The diagnostic segment also serves to steer the beam into the RFQ or terminate the beam as needed. Even small, sub-millimeter misalignment can prevent efficient injection of the beam into the RFQ. In order to ensure that the beam enters the RFQ at both the correct position and angle, two magnetic steerers and a solenoid magnet are placed around the beamline. These are controlled independently by an external power source to produce tunable dipole fields which optimize beam entry into the RFQ. The optimal fields were calculated using COMSOL and beam dynamics simulated in WARP. The effect of the steerers is demonstrated in Fig. 2.6.
In order to control the amount of power delivered from the ion source, a beam chopper is built into the cross of the diagnostic segment to allow for a duty factor in an otherwise DC beam. The beam chopper is built into the cross in the diagnostic section. The chopper comprises a deflection plate, followed by a copper ring for particle dumping press-fitted into a steel beam pipe which in turn is surrounded by a copper water cooling jacket. The deflection plate can be put at a high potential relative to the cross, which is grounded, diverting the beam such that it terminates on the copper ring. See Fig. 2.8. This method permits controlling the average beam power, without changing the charger per bunch resulting from the RFQ. This is required to maintain vortex motion in the cyclotron.
A 12 mA beam of 15 keV particles fully terminating on the beam stop corresponds to $\sim$180 W power delivered. We need to ensure that the dump is capable of dissipating this amount of power. The system was modeled by assuming a water cooling line at room temperature. To be conservative, the simulation was run using four times the actual anticipated beam power, i.e. 720 W. The beam dump never rose above 360 K in our simulation, significantly below the melting point of copper. The COMSOL simulation is shown in Fig. 2.8
Figure 2.7: CAD model of the LEBT diagnostic segment with beam simulation (red) overlaid. If the copper electrode (top) is held at adequate potential the beam is pushed off its center course and terminates on the copper ring. From Ref. [waites_high_2022].
Figure 2.8: COMSOL simulation of the beam terminating on the copper aperture surrounded by the steel beampipe, with copper water cooling channel on the exterior.From Ref. [waites_high_2022].